# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 2.3. Parabolic induction and simple modules
Then this element does not depend on the choice of orientation $o$, and defines an element of the centre of ${\mathcal{H}}$, c.f. [32, Theorem 5.1]. The length of $\tilde{\lambda}\in\mathcal{O}$ is independent of the choice of representative $\tilde{\lambda}$, and we denote it by $\ell(\mathcal{O})$.
@@
Note that the notions of supersingular and uniformly supersingular coincide for ${\mathcal{H}}$-modules of finite length, and more generally finitely generated modules, and hence in particular for simple modules.
@@
Given a supersingular triple $(P,\mathfrak{n},Q)$, we define $I_{\mathcal{H}}(P,\mathfrak{n},Q),CI_{\mathcal{H}}(P,\mathfrak{n},Q)\in{\operatorname{Mod}}({\mathcal{H}})$ as follows. Suppose that $Q$ has Levi decomposition $Q=LV$. Now assume that $Q^{\prime}$ is any parabolic subgroup of $G$ with $P\subseteq Q^{\prime}\subseteq P(\mathfrak{n})$ and say it has Levi decomposition $Q^{\prime}=L^{\prime}V^{\prime}$. Then $\mathfrak{n}$ can be extended uniquely to an ${\mathcal{H}}_{L^{\prime}}$-module $e_{{\mathcal{H}}_{L^{\prime}}}(\mathfrak{n})$ with certain properties, and moreover whenever $P\subseteq Q\subseteq Q^{\prime}\subseteq P(\mathfrak{n})$ there is a natural injective ${\mathcal{H}}$-linear map ${\operatorname{Ind}}_{{\mathcal{H}}_{L^{\prime}}}^{\mathcal{H}}(e_{{\mathcal{H}}_{L^{\prime}}}(\mathfrak{n}))\to{\operatorname{Ind}}_{{\mathcal{H}}_{L}}^{\mathcal{H}}(e_{{\mathcal{H}}_{L}}(\mathfrak{n}))$ (cf. [4, Corollary 3.9 & Proposition 4.5]). We may then define the modules
$I_{\mathcal{H}}(P,\mathfrak{n},Q):={\operatorname{coker}}\left(\bigoplus_{Q\subseteq Q^{\prime}\subseteq P(\mathfrak{n})}{\operatorname{Ind}}_{{\mathcal{H}}_{L^{\prime}}}^{\mathcal{H}}(e_{{\mathcal{H}}_{L^{\prime}}}(\mathfrak{n}))\to{\operatorname{Ind}}_{{\mathcal{H}}_{L}}^{\mathcal{H}}(e_{{\mathcal{H}}_{L}}(\mathfrak{n}))\right)$
@@
$CI_{\mathcal{H}}(P,\mathfrak{n},Q):={\operatorname{coker}}\left(\bigoplus_{Q\subseteq Q^{\prime}\subseteq P(\mathfrak{n})}{\operatorname{Coind}}_{{\mathcal{H}}_{L^{\prime}}}^{\mathcal{H}}(e_{{\mathcal{H}}_{L^{\prime}}}(\mathfrak{n}))\to{\operatorname{Coind}}_{{\mathcal{H}}_{L}}^{\mathcal{H}}(e_{{\mathcal{H}}_{L}}(\mathfrak{n}))\right).$
The classification of simple ${\mathcal{H}}$-modules states that $(P,\mathfrak{n},Q)\mapsto I_{\mathcal{H}}(P,\mathfrak{n},Q)$ and $(P,\mathfrak{n},Q)\mapsto CI_{\mathcal{H}}(P,\mathfrak{n},Q)$ define two bijections between the set of equivalence classes of standard simple supersingular triples and the set of isomorphism classes of simple ${\mathcal{H}}$-modules, cf. [1, Theorem 4.22] and [4, Corollary 4.30]. Moreover, the two bijections are related by an ${\mathcal{H}}$-equivariant isomorphism
@@
cf. [4, Corollary 4.24]. Finally, a simple module $I_{\mathcal{H}}(P,\mathfrak{n},Q)$ is supersingular if and only if $P=Q=G$ (cf. [1, Corollary 4.31]), and analogously for $CI_{\mathcal{H}}(P,\mathfrak{n},Q)$.
@@ 2.4. Recollections on abelian model categories
We briefly recall some facts about model structures on abelian categories. For details on the general theory of model categories, we refer the reader to [9, §1] and the references therein. Fix a bicomplete abelian category ${\mathcal{A}}$. We will say that a model structure on ${\mathcal{A}}$ is an abelian model structure if the class of (trivial) fibrations (resp. cofibrations) coincides with the class of epimorphisms (resp. monomorphisms) with (trivially) fibrant (resp. cofibrant) kernel (resp. cokernel). An abelian category equipped with such a model structure will be called an abelian model category.
By work of Hovey, c.f. [15], constructing abelian model structures on ${\mathcal{A}}$ is equivalent to constructing certain cotorsion pairs. Given a pair $({\mathcal{C}},{\mathcal{F}})$ of classes of objects of ${\mathcal{A}}$, consider the pair $(^{\perp}{\mathcal{F}},{\mathcal{C}}^{\perp})$ defined by
@@
Then we say that $({\mathcal{C}},{\mathcal{F}})$ is a cotorsion pair if ${\mathcal{C}}={{}^{\perp}}{\mathcal{F}}$ and ${\mathcal{F}}={\mathcal{C}}^{\perp}$ (c.f. [15, Definition 2.3]). It is said to have enough functorial projectives if for all $X\in{\mathcal{A}}$, there is a functorial epimorphism $f:Y\to X$ with $Y\in{\mathcal{C}}$ such that $\ker(f)\in{\mathcal{F}}$. Dually, it is said to have enough functorial injectives if for all $X\in{\mathcal{A}}$, there is a functorial monomorphism $g:X\to Z$ with $Z\in{\mathcal{F}}$ such that ${\operatorname{coker}}(g)\in{\mathcal{C}}$. When a cotorsion pair has both enough functorial projectives and enough functorial injectives, then it is said to be functorially complete.
@@
In fact, Hovey proved a more general version of the above theorem which applies to a wider class of model structures on ${\mathcal{A}}$, using the notion of a proper class of short exact sequences, and this has been further generalised to the context of exact categories by Gillespie in [11]. In this paper, we will however only need to work with abelian model structures. We will sometimes denote an abelian model structure on ${\mathcal{A}}$ by the triple $({\mathcal{C}},{\mathcal{F}},{\mathcal{W}})$ as in 2.7. Note that the homotopy category ${\operatorname{Ho}}({\mathcal{A}})$ of an abelian model category is canonically equivalent to the quotient ${\mathcal{C}}\cap{\mathcal{F}}/\Gamma$, where $\Gamma={\mathcal{C}}\cap{\mathcal{F}}\cap{\mathcal{W}}$ (c.f. [11, Proposition 4.4(5)]).
@@
As a final piece of notation, if a functor ${\mathsf{F}}:{\mathcal{A}}\to{\mathcal{B}}$ between the underlying categories of two abelian model categories preserves all weak equivalences, it immediately follows from the universal property of localisation that it lifts to a functor between the homotopy categories which we will denote by ${\operatorname{Ho}}({\mathsf{F}}):{\operatorname{Ho}}({\mathcal{A}})\to{\operatorname{Ho}}({\mathcal{B}})$. When $F$ is Quillen we can identify ${\operatorname{Ho}}({\mathsf{F}})$ with the total derived functor of $F$ under mild assumptions as follows.
@@
Suppose that ${\mathcal{A}}$ and ${\mathcal{B}}$ are abelian model categories, and let ${\mathsf{F}}:{\mathcal{A}}\to{\mathcal{B}}$ be a (left or right) Quillen functor which is exact. Then:
@@
    if ${\mathsf{F}}$ preserves trivial objects then it preserves all weak equivalences; and
@@
    if ${\mathsf{F}}$ reflects trivial objects then it reflects all weak equivalences.
Furthermore, with the assumption of (i), the total derived functor of ${\mathsf{F}}$ is naturally isomorphic to ${\operatorname{Ho}}({\mathsf{F}})$.
@@
Suppose that $f$ is a morphism in ${\mathcal{A}}$. For (i), assume first that $f$ is either a trivial fibration or trivial cofibration. Then in particular it is either a monomorphism with trivial cokernel, or an epimorphism with trivial kernel. Under our assumptions then ${\mathsf{F}}f$ is still a monomorphism/epimorphism with trivial cokernel/kernel. It then follows from [15, Lemma 5.8] that ${\mathsf{F}}f$ is a weak equivalence. The case of a general $f$ follows immediately by factorising $f$ as the composite of a trivial cofibration and a trivial fibration.
For (ii), we assume that ${\mathsf{F}}$ is left Quillen (the other case is dual). Factorise $f=gh$ where $h$ is a trivial cofibration and $g$ is a fibration. Then by the 2-out-of-3 property of weak equivalences it suffices to show that if ${\mathsf{F}}g$ is a weak equivalence then so is $g$. Since ${\mathsf{F}}$ is exact, we may apply loc. cit. again to see that ${\mathsf{F}}g$ is a weak equivalence precisely when ${\mathsf{F}}({\operatorname{coker}}(g))$ is trivial. It then follows that $g$ is a trivial fibration since ${\mathsf{F}}$ reflects trivial objects.
@@
An abelian model category ${\mathcal{A}}$ is called hereditary if the class ${\mathcal{C}}$ of cofibrant objects is closed under taking kernels of epimorphisms and if the class ${\mathcal{F}}$ of fibrant objects is closed under taking cokernels of monomorphisms. Such an abelian model category is automatically stable (c.f. [7, Corollary 1.1.15]), meaning that its associated homotopy category ${\operatorname{Ho}}({\mathcal{A}})$ is canonically equipped with the structure of a triangulated category. We quickly recall what this triangulated structure is.
  • The shift functor is given by the suspension functor $\Sigma$ associated with the model structure. For a cofibrant object $X\in{\mathcal{C}}$, choose a cofibration $j_{X}:X\to W$ where $W\in\Gamma={\mathcal{C}}\cap{\mathcal{F}}\cap{\mathcal{W}}$. Then we let $\Sigma(X)$ be defined as ${\operatorname{coker}}(j_{X})$. Since ${\mathcal{A}}$ is hereditary, note that $\Sigma(X)\in{\mathcal{C}}\cap{\mathcal{F}}$ whenever $X\in{\mathcal{C}}\cap{\mathcal{F}}$. This construction doesn’t depend on the choice of $j_{X}$ up to weak equivalence, and hence this defines an endofunctor of ${\operatorname{Ho}}({\mathcal{A}})\equiv{\mathcal{C}}\cap{\mathcal{F}}/\Gamma$ which can be shown to be an auto-equivalence. There is also a dually defined loop functor which is naturally isomorphic to $\Sigma^{-1}$. We note that the loop and suspension functors can be defined more generally for any (not necessarily stable) pointed model category, c.f. [6, Definition 3.1.2], and are then adjoint functors.
@@
There is a dual notion of fiber sequences which may be used to construct the distinguished triangles as well, c.f. [6, Definition 3.4.2, Corollaries 4.4.1 & 4.4.2]. In our context, those are obtained similarly to the above by using short exact sequences of fibrant objects and the loop functor. That these two constructions give the same triangulated structure on the homotopoy category holds in the general context of stable model categories (c.f. loc. cit.) but for hereditary abelian model categories this can be also seen directly by applying [7, Lemma 1.4.4] and its dual.
@@
Theorem 2.10 ([25, Theorem 0.1]).
[…]
The bound on the selfinjective dimension of ${\mathcal{H}}$ is actually sharp (c.f. [19, Remark 8.2]). The theorem applies generally to the algebras ${\mathcal{H}}_{M}$ for any standard Levi subgroup $M$ of $G$.
@@
of projective objects of ${\mathcal{A}}$ for which $A=Z_{0}Y={\operatorname{Ker}}(d_{0})$ and which remains exact upon applying ${\operatorname{Hom}}_{{\mathcal{A}}}(\cdot,P)$ for any projective object $P$ of ${\mathcal{A}}$. When ${\mathcal{A}}={\operatorname{Mod}}(R)$ for a Gorenstein ring $R$, then a module ${\mathfrak{m}}$ is Gorenstein projective as an object of ${\operatorname{Mod}}(R)$ if and only if ${\operatorname{Ext}}^{i}_{R}({\mathfrak{m}},\mathfrak{p})=0$ for all $i\geq 1$ and all projective $R$-modules $\mathfrak{p}$ (cf. [10, Corollary 11.5.3]). This is true if and only if ${\operatorname{Ext}}^{i}_{R}({\mathfrak{m}},\mathfrak{p})=0$ for all $i\geq 1$ and all modules $\mathfrak{p}$ of finite projective dimension. Then ${\mathfrak{m}}$ is Gorenstein projective if and only if there is an acyclic complex $Y$ of projectives with ${\mathfrak{m}}=Z_{0}Y$, c.f. [9, Remark 3.1]. There is also dually a notion of Gorenstein injective object of ${\mathcal{A}}$, and for ${\mathcal{A}}={\operatorname{Mod}}(R)$ these have dual properties to the above (c.f. [10, Definition 10.1.1 & Corollary 11.2.2])). We denote the classes of Gorenstein projective and Gorenstein injective modules over a ring $R$ by $\mathrm{GProj}(R)$ and $\mathrm{GInj}(R)$, respectively.
@@
The identity functor ${\operatorname{id}}:{\operatorname{Mod}}(R)^{\text{GP}}\to{\operatorname{Mod}}(R)^{\text{GI}}$ is left Quillen and in fact a Quillen equivalence (cf. [15, bottom of p.583]). The latter means that if ${\mathfrak{m}},\mathfrak{n}\in{\operatorname{Mod}}(R)$ are such that ${\mathfrak{m}}$ is Gorenstein projective and $\mathfrak{n}$ is Gorenstein injective, then a map between them is a weak equivalence in one model structure if and only if it is a weak equivalence in the other. But this holds because, more generally, the weak equivalences in both model structures are the same. Indeed, that follows from 2.8 as the trivial objects are the same in either model structure. Even without using the notion of Quillen equivalences, we therefore see that the two model structures have equivalent homotopy categories as these are localisations of ${\operatorname{Mod}}(R)$ with respect to the same collection of weak equivalences. We will always denote this homotopy category by ${\operatorname{Ho}}(R)$.
@@
Under our assumptions, both $F^{B}_{A}$ and ${\operatorname{Res}}^{B}_{A}$ are exact. We claim that they preserve trivial objects, i.e. modules of finite projective/injective dimension. Indeed, $F^{B}_{A}$ and ${\operatorname{Res}}^{B}_{A}$ preserve projective and injective modules, respectively, since they are exact and ajoint. It therefore follows by exactness that $F^{B}_{A}$, resp. ${\operatorname{Res}}^{B}_{A}$, preserves finite projective resolutions, resp. finite injective resolutions, and thus we get our claim. From all of this we also obtain that ${\operatorname{Res}}^{B}_{A}$ preserves epimorphisms, i.e. fibrations, and epimorphisms with kernels of finite projective dimension, i.e. trivial fibrations, and hence is right Quillen as required. Finally, the isomorphism on the total derived functors is immediate from 2.8. ∎
@@
For a Gorenstein ring $R$ and writing ${\mathcal{W}}$ for the class of $R$-modules of finite projective dimension, it is shown in [15, Corollary 8.5] that the classes $\mathrm{GProj}(R)\cap{\mathcal{W}}$ and $\mathrm{GInj}(R)\cap{\mathcal{W}}$ are that of projective and injective $R$-modules respectively. We thus see from 2.14 that for ${\mathfrak{m}}$ a Gorenstein projective module and $\mathfrak{n}$ any module, two maps $f,g:{\mathfrak{m}}\to\mathfrak{n}$ are equal in ${\operatorname{Ho}}(R)$ if and only if $f-g$ factors through a projective module. Similarly, if ${\mathfrak{m}}$ is any module and if $\mathfrak{n}$ is Gorenstein injective, then two maps $f,g:{\mathfrak{m}}\to\mathfrak{n}$ are equal in ${\operatorname{Ho}}(R)$ if and only if $f-g$ factors through an injective module.

### v4

@@ 2.3. Parabolic induction and simple modules
Then this element does not depend on the choice of orientation $o$, and defines an element of the centre of ${\mathcal{H}}$, cf. [35, Theorem 5.1]. The length of $\tilde{\lambda}\in\mathcal{O}$ is independent of the choice of representative $\tilde{\lambda}$, and we denote it by $\ell(\mathcal{O})$.
@@
Note that the notions of supersingular and uniformly supersingular coincide for ${\mathcal{H}}$-modules of finite length, or more generally for finitely generated modules, and hence in particular for simple modules.
@@
Given a supersingular triple $(P,\mathfrak{n},Q)$, we define $I_{\mathcal{H}}(P,\mathfrak{n},Q),CI_{\mathcal{H}}(P,\mathfrak{n},Q)\in{\operatorname{Mod}}({\mathcal{H}})$ as follows. Suppose that $Q$ has Levi decomposition $Q=LV$. Now assume that $Q^{\prime}$ is any parabolic subgroup of $G$ with $P\subseteq Q^{\prime}\subseteq P(\mathfrak{n})$ and say it has Levi decomposition $Q^{\prime}=L^{\prime}V^{\prime}$. Then $\mathfrak{n}$ can be extended uniquely to an ${\mathcal{H}}_{L^{\prime}}$-module $e_{{\mathcal{H}}_{L^{\prime}}}(\mathfrak{n})$ with certain properties, and moreover whenever $P\subseteq Q\subseteq Q^{\prime}\subseteq P(\mathfrak{n})$ there is a natural injective ${\mathcal{H}}$-linear map ${\operatorname{Ind}}_{{\mathcal{H}}_{L^{\prime}}}^{\mathcal{H}}(e_{{\mathcal{H}}_{L^{\prime}}}(\mathfrak{n}))\to{\operatorname{Ind}}_{{\mathcal{H}}_{L}}^{\mathcal{H}}(e_{{\mathcal{H}}_{L}}(\mathfrak{n}))$ (cf. [4, Corollary 3.9 & Proposition 4.5]). There is also an analogous map for coinduction, cf. the discussion preceding [1, Theorem 4.22]. We may thus define the modules
(2.8) $I_{\mathcal{H}}(P,\mathfrak{n},Q):={\operatorname{coker}}\left(\bigoplus_{Q\subsetneq Q^{\prime}\subseteq P(\mathfrak{n})}{\operatorname{Ind}}_{{\mathcal{H}}_{L^{\prime}}}^{\mathcal{H}}(e_{{\mathcal{H}}_{L^{\prime}}}(\mathfrak{n}))\to{\operatorname{Ind}}_{{\mathcal{H}}_{L}}^{\mathcal{H}}(e_{{\mathcal{H}}_{L}}(\mathfrak{n}))\right)$
@@
(2.9) $CI_{\mathcal{H}}(P,\mathfrak{n},Q):={\operatorname{coker}}\left(\bigoplus_{Q\subsetneq Q^{\prime}\subseteq P(\mathfrak{n})}{\operatorname{Coind}}_{{\mathcal{H}}_{L^{\prime}}}^{\mathcal{H}}(e_{{\mathcal{H}}_{L^{\prime}}}(\mathfrak{n}))\to{\operatorname{Coind}}_{{\mathcal{H}}_{L}}^{\mathcal{H}}(e_{{\mathcal{H}}_{L}}(\mathfrak{n}))\right).$
The classification of simple ${\mathcal{H}}$-modules states that $(P,\mathfrak{n},Q)\mapsto I_{\mathcal{H}}(P,\mathfrak{n},Q)$ and $(P,\mathfrak{n},Q)\mapsto CI_{\mathcal{H}}(P,\mathfrak{n},Q)$ define two bijections between the set of equivalence classes of standard simple supersingular triples and the set of isomorphism classes of simple ${\mathcal{H}}$-modules, cf. loc. cit. and [4, Corollary 4.30]. Moreover, the two bijections are related by an ${\mathcal{H}}$-equivariant isomorphism
@@
cf. [4, Corollary 4.24]. Finally, a simple module $CI_{\mathcal{H}}(P,\mathfrak{n},Q)$ is supersingular if and only if $P=Q=G$ (cf. [1, Corollary 4.31]), and analogously for $I_{\mathcal{H}}(P,\mathfrak{n},Q)$.
@@ 2.4. Recollections on abelian model categories
We briefly recall some facts about model structures on abelian categories. For details on the general theory of model categories, we refer the reader to [9, §1] and the references therein. Fix a bicomplete abelian category ${\mathcal{A}}$. We will say that a model structure on ${\mathcal{A}}$ is an abelian model structure if the class of (trivial) fibrations, resp. (trivial) cofibrations, coincides with the class of epimorphisms with (trivially) fibrant kernel, resp. monomorphisms with (trivially) cofibrant cokernel. An abelian category equipped with such a model structure will be called an abelian model category.
By work of Hovey, cf. [17], constructing abelian model structures on ${\mathcal{A}}$ is equivalent to constructing certain cotorsion pairs. Given a pair $({\mathcal{C}},{\mathcal{F}})$ of classes of objects of ${\mathcal{A}}$, consider the pair $(^{\perp}{\mathcal{F}},{\mathcal{C}}^{\perp})$ defined by
@@
Then we say that $({\mathcal{C}},{\mathcal{F}})$ is a cotorsion pair if ${\mathcal{C}}={{}^{\perp}}{\mathcal{F}}$ and ${\mathcal{F}}={\mathcal{C}}^{\perp}$ (cf. [17, Definition 2.3]). It is said to have enough functorial projectives if for all $X\in{\mathcal{A}}$, there is a functorial epimorphism $f:Y\to X$ with $Y\in{\mathcal{C}}$ such that $\ker(f)\in{\mathcal{F}}$. Dually, it is said to have enough functorial injectives if for all $X\in{\mathcal{A}}$, there is a functorial monomorphism $g:X\to Z$ with $Z\in{\mathcal{F}}$ such that ${\operatorname{coker}}(g)\in{\mathcal{C}}$. When a cotorsion pair has both enough functorial projectives and enough functorial injectives, then it is said to be functorially complete.
@@
In fact, Hovey proved a more general version of the above theorem which applies to a wider class of model structures on ${\mathcal{A}}$, using the notion of a proper class of short exact sequences, and this has been further generalised to the context of exact categories by Gillespie in [11]. In this paper, we will however only need to work with abelian model structures. We will sometimes denote an abelian model structure on ${\mathcal{A}}$ by the triple $({\mathcal{C}},{\mathcal{F}},{\mathcal{W}})$ as in 2.7. Note that the homotopy category ${\operatorname{Ho}}({\mathcal{A}})$ of an abelian model category is canonically equivalent to the quotient ${\mathcal{C}}\cap{\mathcal{F}}/\Gamma$, where $\Gamma={\mathcal{C}}\cap{\mathcal{F}}\cap{\mathcal{W}}$ (cf. [11, Proposition 4.4(5)]).
@@
As a final piece of notation, if a functor ${\mathsf{F}}:{\mathcal{A}}\to{\mathcal{B}}$ between the underlying categories of two abelian model categories preserves all weak equivalences, it then naturally induces a functor ${\operatorname{Ho}}({\mathsf{F}}):{\operatorname{Ho}}({\mathcal{A}})\to{\operatorname{Ho}}({\mathcal{B}})$ given by ${\operatorname{Ho}}({\mathsf{F}})(X):={\mathsf{F}}(X)$ for $X\in{\operatorname{Ho}}({\mathcal{A}})$. When ${\mathsf{F}}$ is Quillen we can identify ${\operatorname{Ho}}({\mathsf{F}})$ with the total derived functor of ${\mathsf{F}}$ under mild assumptions as follows.
@@
Suppose that ${\mathcal{A}}$ and ${\mathcal{B}}$ are abelian model categories, and let ${\mathsf{F}}:{\mathcal{A}}\to{\mathcal{B}}$ be an exact functor between the underlying abelian categories.
@@
    If ${\mathsf{F}}$ preserves trivial objects then it preserves all weak equivalences.
@@
    Assume that ${\mathsf{F}}$ is a (left or right) Quillen functor. If ${\mathsf{F}}$ reflects trivial objects then it reflects all weak equivalences.
Furthermore, with the assumption of (i) and if ${\mathsf{F}}$ is Quillen, then the total derived functor of ${\mathsf{F}}$ is naturally isomorphic to ${\operatorname{Ho}}({\mathsf{F}})$.
@@
Suppose that $f$ is a morphism in ${\mathcal{A}}$. For (i), assume first that $f$ is either a trivial cofibration or trivial fibration. Then in particular it is either a monomorphism with trivial cokernel, or an epimorphism with trivial kernel. Under our assumptions then ${\mathsf{F}}f$ is still a monomorphism/epimorphism with trivial cokernel/kernel. It then follows from [17, Lemma 5.8] that ${\mathsf{F}}f$ is a weak equivalence. The case of a general $f$ follows immediately by factorising $f$ as the composite of a trivial cofibration and a trivial fibration.
For (ii), we assume that ${\mathsf{F}}$ is left Quillen (the other case is dual). Factorise $f=gh$ where $h$ is a trivial cofibration and $g$ is a fibration. Then by the 2-out-of-3 property of weak equivalences it suffices to show that if ${\mathsf{F}}g$ is a weak equivalence then so is $g$. Since ${\mathsf{F}}$ is exact, we may apply loc. cit. again to see that ${\mathsf{F}}g$ is a weak equivalence precisely when ${\mathsf{F}}(\ker(g))$ is trivial. It then follows that $g$ is a trivial fibration since ${\mathsf{F}}$ reflects trivial objects.
@@
An abelian model category ${\mathcal{A}}$ is called hereditary if the class ${\mathcal{C}}$ of cofibrant objects is closed under taking kernels of epimorphisms and if the class ${\mathcal{F}}$ of fibrant objects is closed under taking cokernels of monomorphisms. Such an abelian model category is automatically stable (cf. [7, Corollary 1.1.15]), meaning that its associated homotopy category ${\operatorname{Ho}}({\mathcal{A}})$ is canonically equipped with the structure of a triangulated category. We quickly recall what this triangulated structure is.
  • The shift functor is given by the suspension functor $\Sigma$ associated with the model structure. For a cofibrant object $X\in{\mathcal{C}}$, we may choose from the factorisations in the model structure a cofibration $j_{X}:X\to W$ where $W$ is trivially fibrant, and in fact $W\in\Gamma={\mathcal{C}}\cap{\mathcal{F}}\cap{\mathcal{W}}$ since cofibrations are closed under compositions. Then we let $\Sigma(X)$ be defined as ${\operatorname{coker}}(j_{X})$. Since ${\mathcal{A}}$ is hereditary, note that $\Sigma(X)\in{\mathcal{C}}\cap{\mathcal{F}}$ whenever $X\in{\mathcal{C}}\cap{\mathcal{F}}$. This construction doesn’t depend on the choice of $j_{X}$ up to weak equivalence, and hence this defines an endofunctor of ${\operatorname{Ho}}({\mathcal{A}})\simeq{\mathcal{C}}\cap{\mathcal{F}}/\Gamma$ which can be shown to be an auto-equivalence. There is also a dually defined loop functor which is naturally isomorphic to $\Sigma^{-1}$. We note that the loop and suspension functors can be defined more generally for any (not necessarily stable) pointed model category, cf. [6, Definition 3.1.2], and are then adjoint functors.
@@
There is a dual notion of fiber sequences which may be used to construct the distinguished triangles as well, cf. [6, Definition 3.4.2, Corollaries 4.4.1 & 4.4.2]. In our context, those are obtained similarly to the above by using short exact sequences of fibrant objects and the loop functor. That these two constructions give the same triangulated structure on the homotopy category holds in the general context of stable model categories (cf. loc. cit.) but for hereditary abelian model categories this can be also seen directly by applying [7, Lemma 1.4.4] and its dual.
@@
Theorem 2.10 ([27, Theorem 0.1]).
The pro-$p$ Iwahori-Hecke algebra ${\mathcal{H}}$ is $n$-Gorenstein where $n\leq r_{\text{ss}}+r_{Z}$.
The bound on the selfinjective dimension of ${\mathcal{H}}$ is actually sharp (cf. [21, Remark 8.2]). The theorem applies generally to the algebras ${\mathcal{H}}_{M}$ for any standard Levi subgroup $M$ of $G$.
@@
of projective objects of ${\mathcal{A}}$ for which $A=Z_{0}Y={\operatorname{Ker}}(d_{0})$ and which remains exact upon applying ${\operatorname{Hom}}_{{\mathcal{A}}}(\cdot,P)$ for any projective object $P$ of ${\mathcal{A}}$. When ${\mathcal{A}}={\operatorname{Mod}}(R)$ for a Gorenstein ring $R$, then a module ${\mathfrak{m}}$ is Gorenstein projective as an object of ${\operatorname{Mod}}(R)$ if and only if ${\operatorname{Ext}}^{i}_{R}({\mathfrak{m}},\mathfrak{p})=0$ for all $i\geq 1$ and all modules $\mathfrak{p}$ of finite projective dimension (cf. [10, Corollary 11.5.3]). Then ${\mathfrak{m}}$ is Gorenstein projective if and only if there is an acyclic complex $Y$ of projectives with ${\mathfrak{m}}=Z_{0}Y$, cf. [9, Remark 3.1]. There is also dually a notion of Gorenstein injective object of ${\mathcal{A}}$, and for ${\mathcal{A}}={\operatorname{Mod}}(R)$ these have dual properties to the above (cf. [10, Definition 10.1.1 & Corollary 11.2.2])). We denote the classes of Gorenstein projective and Gorenstein injective modules over a ring $R$ by $\mathrm{GProj}(R)$ and $\mathrm{GInj}(R)$, respectively.
@@
The identity adjunction
${\operatorname{id}}:{\operatorname{Mod}}(R)^{\text{GP}}\rightleftarrows{\operatorname{Mod}}(R)^{\text{GI}}:{\operatorname{id}}$
is Quillen and in fact a Quillen equivalence (cf. [17, bottom of p.583]). The latter means that if ${\mathfrak{m}},\mathfrak{n}\in{\operatorname{Mod}}(R)$ are such that ${\mathfrak{m}}$ is Gorenstein projective and $\mathfrak{n}$ is Gorenstein injective, then a map between them is a weak equivalence in one model structure if and only if it is a weak equivalence in the other. But this holds because, more generally, the weak equivalences in both model structures are the same. Indeed, that follows from 2.8 as the trivial objects are the same in either model structure.
Even without using the notion of Quillen equivalences, we therefore see that the two model structures have the same homotopy category, namely the abstract localisation ${\operatorname{Ho}}(R):={\operatorname{Mod}}(R)[(\text{w.e.})^{-1}]$ as defined in [16, Definition 1.2.1] (where ‘w.e.’ stands for ‘weak equivalences’). Our two model structures provide two categories, namely $\mathrm{GProj}(R)/\mathrm{Proj}(R)$ and $\mathrm{GInj}(R)/\mathrm{Inj}(R)$, which are equivalent to it and in particular which describe the ${\operatorname{Hom}}$ sets explicitly. Similarly, the two model structures give two different collections of representatives for the isomorphism classes of the same distinguished triangles in ${\operatorname{Ho}}(R)$, and similarly give two different (explicit) constructions for the same shift functor $\Sigma$ (see also 3.12 later). In this paper we will use both of these model structures, depending on the situation.
@@
Under our assumptions, both $F^{B}_{A}$ and ${\operatorname{Res}}^{B}_{A}$ are exact. We claim that they preserve trivial objects, i.e. modules of finite projective/injective dimension. Indeed, $F^{B}_{A}$ and ${\operatorname{Res}}^{B}_{A}$ preserve projective and injective modules, respectively, since they are an adjoint pair and exact. It therefore follows by exactness that $F^{B}_{A}$, resp. ${\operatorname{Res}}^{B}_{A}$, preserves finite projective resolutions, resp. finite injective resolutions, and thus we get our claim. From all of this we also obtain that ${\operatorname{Res}}^{B}_{A}$ preserves epimorphisms, i.e. fibrations, and epimorphisms with kernels of finite projective dimension, i.e. trivial fibrations, and hence is right Quillen as required. Finally, the isomorphism on the total derived functors is immediate from 2.8. ∎
@@
For a Gorenstein ring $R$ and writing ${\mathcal{W}}$ for the class of $R$-modules of finite projective dimension, it is shown in [17, Corollary 8.5] that the classes $\mathrm{GProj}(R)\cap{\mathcal{W}}$ and $\mathrm{GInj}(R)\cap{\mathcal{W}}$ are that of projective and injective $R$-modules respectively. We thus see from 2.14 that for ${\mathfrak{m}}$ a Gorenstein projective module and $\mathfrak{n}$ any module, two maps $f,g:{\mathfrak{m}}\to\mathfrak{n}$ are equal in ${\operatorname{Ho}}(R)$ if and only if $f-g$ factors through a projective module. Similarly, if ${\mathfrak{m}}$ is any module and if $\mathfrak{n}$ is Gorenstein injective, then two maps $f,g:{\mathfrak{m}}\to\mathfrak{n}$ are equal in ${\operatorname{Ho}}(R)$ if and only if $f-g$ factors through an injective module.